# BanTH: A Multi-label Hate Speech Detection Dataset for Transliterated Bangla

### v3

@@ 4 Experimental Result
We conduct experiments on fine-tuning LMs along with utilizing prompt-based techniques. Our experiment setup for finetuning LMs is described in Appendix D and the prompts are reported in Appendix E. Table 5 shows the benchmarking of various models on our dataset for both binary and multi-label classification tasks.
@@ 4.1 Fine-Tuning LMs Baseline
For the binary classification, among the fine-tuned models, TB-mBERT achieves the highest accuracy (82.57%) and Macro-F1 score (77.36%). The best-performing BERT variant is TB-BanglaBERT, with an accuracy of 81.61% and Macro-F1 of 77.12%. In multi-label classification, for Macro-F1, TB-BERT leads with 30.17%, followed closely by XLM-RoBERTa (28.24%). In subset accuracy, TB-mBERT performs best (54.71%), with TB-BERT close behind (54.19%). Hamming Loss is lowest (best) for TB-BanglaBERT (7.26). From the experiments, our proposed TB-encoders beat the other LMs for both binary and multi-label classification.
@@ 4.2 Prompt Based LLM Performance
Among prompt-based LLMs, GPT-4o + Few-shot performed the best for multi-label classification, achieving the highest Macro-F1 (39.53%), Subset Accuracy (26.76%), and lowest Hamming Loss (14.16%). In binary classification, GPT 4o+Exp+Few-shot has the highest Macro-F1 (70.70%). In terms of accuracy, our Translation Prompting approach gains the highest result (74.87%). If we consider zero-shot prompting techniques, the proposed approach consistently outperforms all other prompt variations.
Our extensive experiments reveal that fine-tuned models outperform all prompt-based results for both binary and multi-label classification with a good margin (~8% improvement over prompting). The few-shot prompting strategy enhances model performance compared to the zero-shot approach, although this is not true for all cases. Additionally, GPT-4 consistently outperforms GPT-3.5 across all experiments and prompting variations, yielding improvements of 8-10%. We also find that generating explanations can occasionally boost performance, but it may also lead to worse results in some instances. Our proposed prompting techniques achieve the best performance in zero-shot settings. Overall, our proposed TB encoders demonstrate superior performance, surpassing all fine-tuned LMs and prompt-based results.
@@ 5 Error Analysis
Figure 4 illustrates per-label accuracy with the best performing model, TB-mBert. From the fig, it is evident that the classification accuracy for all the labels except Personal Offense label. Figure 3 indicates that most hate speech labels are categorized under Personal Offense. Consequently, the subset accuracy of the multi-label classification, as presented in Table 5, is lower than the per-label accuracy. This suggests that the model struggles to effectively classify samples of Personal Offense label. As illustrated by Fig. 5, the minimum accuracy for binary and multi-label classification is at least 80% and 50%, respectively, with Entertainment videos holding the highest accuracy in both the classifications, showing higher subset accuracy in multi-label classification (79.5%).
It is evident from Tab. 6 that the number of Personal Offense in Entertainment video type is comparable to that of the other multi-labels of hate, whereas in the other two categories, News & Politics and People & Blogs, the number of Personal Offense hate is noticeably high. The overall improved accuracy in Entertainment category aligns with the result produced in Fig. 4.
The results in Tab. 5 show that our Translation-based models, like GPT 4o Bangla/English Translation with explanations, perform nearly as well as top prompt-based models. As observed in Tab. 7, LLMs can naturally translate Bangla transliterated text into English, even without explicit instructions. However, when asked to translate, the LLM tends to give more literal translations. We also observe that poor translations often lead to incorrect classifications, especially with regional dialects or terms lacking direct English equivalents. English translations generally outperform Bangla ones. When not prompted to translate, the LLM focuses on identifying hate-related terms, which works for explicit hate speech but struggles with subtler forms of offensive content, leading to misclassification.
@@ 6 Related Work
Bangla Hate Speech. As a low-resource language, Bangla has limited work on hate speech detection Romim et al. (2022); Das et al. (2022a) but has made significant progress on other hate-related tasks involving abusive content Hussain et al. (2018); Jahan et al. (2022a), cyberbullying Emon et al. (2022), sexism Jahan et al. (2023), and violence Fahim (2023). The multi-label variants are generally categorized based on religion, politics, and gender Sharif et al. (2022). Key areas of multi-labeling tasks include aggression Hossain et al. (2023), cyberbullying Saifuddin et al. (2023), hate speech Shakil (2022), and toxic content Belal et al. (2023).
@@ 7 Conclusion
The BanTH dataset contains 27167 non-hate samples out of 37350, i.e. a dummy classifier predicting all samples as non-hate achieves an accuracy of 72.73%. This undermines the performance of a few methods, notably, GPT-3.5 performs worse than this baseline. We also recognize that our work does not cover all possible hate classes and may not adequately represent other regionally relevant hate classes, e.g. hate based on socio-economic conditions.
@@
The use of comments scraped from the YouTube videos complies with the YouTube API’s terms of services444https://developers.google.com/youtube/terms/api-services-terms-of-service. All forms of Personal Identification Information (PII) have been removed from the dataset to prevent privacy violations. We ensured fair distribution of annotation workload. The hired annotators and domain experts were compensated on an hourly basis at a rate above the industry standard.
@@ C.1 Formal Description of Target Groups
C.2 Category-wise Multi-label Distribution
| Multi-labels | #News & Politics | #People & Blogs | #Entertainment |
| Political | 1005 | 174 | 21 |
| Religious | 109 | 30 | 0 |
| Gender | 151 | 38 | 4 |
| Personal Offense | 5883 | 784 | 297 |
| Abusive/Violence | 2491 | 276 | 121 |
| Origin | 201 | 11 | 2 |
| Body Shaming | 56 | 29 | 1 |
The distribution of multi-labels within the BanTH dataset in Table 6 reveals notable variances in the prevalence of various types of hate speech across video categories, including News & Politics, People & Blogs, and Entertainment. Table 1 shows that the News & Politics video category holds a major share in the BanTH dataset, so it is less surprising that the number of hate samples of each label will be greater in that. The Personal Offense and Body Shaming labels are the most and least prominent in the News & Politics category, with 5883 and 56 samples respectively. In fact, the Personal Offense label is prominent in every video category.
C.3 Wordcloud
From Figure 6 and the corresponding English translations in Figure 7, we can visualize the mostly mentioned words for each label in the BanTH dataset. For Political label, people often referred to Hasina (ex-prime minister of Bangladesh), bnp (a major political party in Bangladesh), league/lig (wing of a political party in Bangladesh), dalal (broker), police (a law enforcement force), etc. As we have covered transliterated Bangla YouTube comments of July 2024, the BanTH dataset includes the public responses related to the political unrest of Bangladesh during that period. In Religious labeled sample texts, people often mention Muslim, Hindu nastik (atheist), Allah (the one and only God in Islam), etc. Gender related hate comments are mostly related to females. In Personal Offense case, commenters address dalal (broker) and police (a law enforcement force) contemptuously using tui/tor. The same case is evident in Abusive/Violence labeled samples because the police showed violence during the student protest of July 2024. As transliterated Bangla is used by mainly Indian and Bangladeshi people, Origin related hate is often directed towards Bangladeshi and Indian people. For Body Shaming classified samples, we see people refer to others as mota (fat) and takla (bald) to body shame and it is directed mostly towards females as apu (sister) is one of the most used words. Although the BanTH dataset is concentrated on these types of hateful words, it may not be the case when the source is different.
@@ Appendix D Experimental Setup
For Language Models (LMs) baseline experiments, we fine-tuned a range of pre-trained LMs to establish baselines for binary and multi-label classification on the BanTH dataset. The models include multilingual architectures like XLM-RoBERTaConneau et al. (2019) and mBERTDevlin et al. (2018), as well as language-specific models such as BanglaBERTBhattacharjee et al. (2022) and BanglishBERTBhattacharjee et al. (2022), which are designed for Bangla and code-mixed transliterated Bangla text, respectively. We also employed BanglaHateBERTJahan et al. (2022b), specifically optimized for detecting hate speech in Bangla, along with IndicBERTKakwani et al. (2020) and MuRILKhanuja et al. (2021), tailored for Indian languages. Additionally, we included models like CharBERTMa et al. (2020), and VAC-BERTBhattacharyya et al. (2023) to capture diverse model architectures, including lightweight models and character-level representations.
@@ D.1 Performance Metrics
Appendix E Prompts for Classification on BanTH dataset

### v4

@@ 4 Experimental Result
The experimental setup is reported in App. D. We evaluate fine-tuning vanilla pre-trained and further pre-trained LMs and different prompting strategies on LLMs. The evaluation metric details have also been provided in App. E.
@@ 4.1 Fine-Tuning LMs Baseline
For the binary classification, TB-mBERT achieves the best performance across all metrics. In multi-label classification, TB-BERT outperforms the other methods based on macro-F1 and hamming loss but is outperformed by TB-mBERT based on subset accuracy. Overall, the TB encoders have better performance across all metrics in both binary and multi-label classification tasks.
@@ 4.2 Prompt Based LLM Performance
Among prompt-based LLMs, GPT-4o + Few-shot performed the best for multi-label classification, achieving the highest macro-F1 (39.53%), subset accuracy (26.76%), and lowest hamming loss (14.16%). In binary classification, GPT 4o+Exp+Few-shot has the highest macro-F1 (70.70%). In zero-shot settings, our translation prompting strategy achieves the best performance across all metrics by a good margin (~8% improvement over vanilla prompting). Additionally, GPT-4 consistently outperforms GPT-3.5 across all experiments and prompting variations, yielding improvements of 8-10%. We also find that generating explanations can occasionally boost performance, but may also lead to worse results.
@@ 5 Error Analysis
5.1 High Error in Personal Offense
Figure 4 illustrates per-label accuracy with the best performing model, TB-mBERT. Most hate samples are labeled under Personal Offense (Tab. 1). Interestingly, Personal Offense has the lowest macro-F1 score for mBERT (Fig. 4). This is due to Personal Offense exhibiting higher variations of transliterated text (Fig. 6,7), making it difficult to attribute specific words that signify this category. The higher co-occurrence of this category with other categories might also contribute to the higher error.
5.2 Representative Words
Analyzing the word cloud in Fig. 6,7, we observe that certain words representative of their respective hate categories frequently occurs in the hate categories with higher performance. For instance, Muslim frequently appears in "Religious" hate, mohila (woman) and meye (girl) in "Gender" hate, and so on. In contrast, the frequently occurring words for "Personal Offense" appear to be generic terms, e.g. dalal (broker), police, and tui (you). Consequently, it becomes more difficult for language models to classify a word as "Personal Offense" due to the lack of representative words.
5.3 Translation by LLMs
As observed in Tab. 8, LLMs can naturally translate Bangla transliterated text into English, even without explicit instructions. However, when explicitly instructed to translate, the LLM tends to give more literal translations. We also observe that erroneous translations often lead to incorrect classifications, especially with regional dialects or terms lacking direct English equivalents. English translations are generally observed to outperform Bangla ones. When not prompted to translate, the LLM focuses on identifying hate-related terms, which works for explicit hate speech but struggles with subtler forms of offensive content, leading to misclassification.
@@ 6 Related Work
Bangla Hate Speech. As a low-resource language, Bangla has limited work on hate speech detection Romim et al. (2022); Das et al. (2022a) but has made significant progress on other hate-related tasks involving abusive content Hussain et al. (2018); Jahan et al. (2022b), cyberbullying Emon et al. (2022), sexism Jahan et al. (2023), and violence Fahim (2023). The multi-label variants are generally categorized based on religion, politics, and gender Sharif et al. (2022). Key areas of multi-labeling tasks include aggression Hossain et al. (2023), cyberbullying Saifuddin et al. (2023), hate speech Shakil (2022); Hasan et al. (2024), and toxic content Belal et al. (2023).
@@ 7 Conclusion
The BanTH dataset contains 27167 non-hate samples out of 37350, i.e. a dummy classifier predicting all samples as non-hate achieves an accuracy of 72.73%. This undermines the performance of a few methods, notably, GPT-3.5 performs worse than this baseline.
Our work does not cover all possible hate classes and may not adequately represent other regionally relevant categories, e.g. hate based on socio-economic conditions. As the dataset has been sourced from YouTube comments, there are insufficient instances of such categories. Adding a low proportion of these samples will create more imbalance in multi-label classification.
We emphasize that the inter-annotator score is comparatively lower than in other classification tasks, primarily due to the inherently subjective nature of hate speech and the variations in transliteration. The expert agreements have also been found lower than the normal annotators, which can be attributed to experts having higher levels of disagreement in defining hate categories due to their deeper understanding of the gray areas in interpretation. In contrast, the simpler and straightforward thinking can be attributed to the higher agreement scores between general annotators.
@@
The use of comments scraped from the YouTube videos complies with the YouTube API’s terms of services333https://developers.google.com/youtube/terms/api-services-terms-of-service. All forms of Personal Identification Information (PII) have been removed from the dataset to prevent privacy violations. We ensured fair distribution of annotation workload. The hired annotators and domain experts were compensated on an hourly basis at a rate above the industry standard.
Acknowledgments
We sincerely appreciate the generous support of our project sponsor, Penta Global Limited, Bangladesh.
@@ C.1 Formal Description of Target Groups
C.2 Wordcloud
From Figure 6 and the corresponding English translations in Figure 7, we can visualize the mostly mentioned words for each label in the BanTH dataset. For Political label, people often referred to Hasina (ex-prime minister of Bangladesh), bnp (a major political party in Bangladesh), league/lig (wing of a political party in Bangladesh), dalal (broker), police (a law enforcement force), etc. As we have covered transliterated Bangla YouTube comments of July 2024, the BanTH dataset includes the public responses related to the political unrest of Bangladesh during that period.
In Religious labeled sample texts, people often mention Muslim, Hindu nastik (atheist), Allah (the one and only God in Islam), etc. Gender related hate comments are mostly related to females. In Personal Offense case, commenters address dalal (broker) and police (a law enforcement force) contemptuously using tui/tor. The same case is evident in Abusive/Violence labeled samples because the police showed violence during the student protest of July 2024.
As transliterated Bangla is used by mainly Indian and Bangladeshi people, Origin related hate is often directed towards Bangladeshi and Indian people. For Body Shaming classified samples, we see people refer to others as mota (fat) and takla (bald) to body shame and it is directed mostly towards females as apu (sister) is one of the most used words. Although the BanTH dataset is concentrated on these types of hateful words, it may not be the case when the source is different.
@@ Appendix D Experimental Setup
For Language Models (LMs) baseline experiments, we fine-tuned a range of pre-trained LMs to establish baselines for binary and multi-label classification on the BanTH dataset. The models include multilingual architectures like XLM-RoBERTaConneau et al. (2019) and mBERTDevlin et al. (2018), as well as language-specific models such as BanglaBERTBhattacharjee et al. (2022) and BanglishBERTBhattacharjee et al. (2022), which are designed for Bangla and code-mixed transliterated Bangla text, respectively. We also employed BanglaHateBERTJahan et al. (2022a), specifically optimized for detecting hate speech in Bangla, along with IndicBERTKakwani et al. (2020) and MuRILKhanuja et al. (2021), tailored for Indian languages. Additionally, we included models like CharBERTMa et al. (2020), and VAC-BERTBhattacharyya et al. (2023) to capture diverse model architectures, including lightweight models and character-level representations.
@@ Appendix E Performance Metrics
Appendix F Prompts for Classification on BanTH dataset